OpenStudy (anonymous):

What is the value of s in the equation 4(2s − 1) = 7s + 12? Could I please get an explanation as well?

OpenStudy (anonymous):

My answer choices are A. 12 B. 16 C. 32 D. 34

OpenStudy (anonymous):

Geteasysolution.com? the set up looks the same lol. I tried that before coming here but it couldn't pull up the equation. Thank you though!

OpenStudy (anonymous):

First expand the left side

OpenStudy (anonymous):

\[8s-4=7s+12\]Then,combine like terms

OpenStudy (anonymous):

Subtract 7s on both sides\[8s-7s-4=7s-7s+12\]\[8s-7s-4=12\]Then,add 4 on both sides\[8s-7s-4+4=12+4\]\[8s-7s=12+4\]

OpenStudy (anonymous):

\[s=?\]solve it @Lemonmeteorite

OpenStudy (anonymous):

Yeah, sorry this site is extremely glitchy on my end.